find the difference of areas if the radius of the bigger circle is 6 cm. The radius of the bigger circle is the diameter of the smaller circle.
note: you dont have to use the value of pi, the answer may be like this: 9π sq. cm

Answer by rfer(16322)

large circle= 3.14*6^2=113.04 sq cm
small circle=3.14*3^2=28.26 sq cm
difference=3.14*27 or 84.78 sq cm